Need help, Wil upvote thank you!

Assume that total output can be computed by using the following formula:

Total output = number of workers x productivity

Assume that there are 120 workers and each worker produces $200 of output.

How fast can output increase if the workforce increases by 5% and productivity increases by 7? Round your answer to 2 decimal places as a percentage.
Answer #1

Total output = Number of workers * Productivity

Total Productivity with 120 worker and productivity = 200

Total Output = 120 * 200 = 24,000

If workforce rises by 5% which means workforce is 120 + 120 * 0.05 = 126 while productivity rise by 7% which means new productivity is 200 + 200 * 0.07 = 214

Total Output would be = 126 * 214 = 26,964

% change in total output = [(26,964 - 24,000) / 24,000] * 100 = 12.35%